Light enters a diamond, reflects internally several times, and then refracts back into air. What is the critical angle for diamond, given the speed of light in diamond is 1.24x108 m/s ?
1
Expert's answer
2020-06-17T09:30:54-0400

First, calculate the refraction index:


"n=\frac{c}{v}=\frac{3\cdot10^8}{1.24\cdot10^8}=2.42."

the critical angle for air as surrounding medium is


"\theta_c=\text{arcsin}\frac{1}{n}=\text{arcsin}\frac{1}{2.42}=24.4°."